OpenStudy (anonymous):

use the fundemental identities to simplify the expression. 1/tan^2x + 1

OpenStudy (anonymous):

tan^2 x + 1 = sec^2 x

OpenStudy (anonymous):

also 1/sec x = cos x

OpenStudy (anonymous):

but it is sec^2x so is the answer cos^2x?

OpenStudy (anonymous):

What you know is that 1/(tan^2x)+1=1/(sin^2x/cos^2x)+1 which is equal to cos^2x/(sin^2x) + 1, write this on a common denominator.

OpenStudy (anonymous):

At that stage my guess is that you'll easily spot something beautiful! :D